# Fantastic Copyrighted Beasts and How (Not) to Generate Them

### v1

@@ 4 Mitigation Strategies
Prompt rewriting is an existing mitigation used in production-level systems such as DALL·E. Specifically, the DALL·E interface contains a prompt-rewriting step that first processes the user’s text input into a format that DALL·E can use to generate images and comply with OpenAI’s policies, such as avoiding copyrighted content. In order to simulate the prompt-rewriting pipeline, we query GPT-4 with the DALL·E’s full system prompt (see full template in §C.6) and the keywords or descriptions to be rewritten.
Example of applying prompt rewriting for ‘Mario’ ‘Create an image of a fictional character inspired by the world of classic video games. He is a middle-aged man of Italian descent, with a robust physique, and typically clad in a red shirt and blue overalls. His most distinctive features include a bushy mustache and a red cap... ’
@@
Negative prompts are often used in deployed diffusion model deployments (Playground AI, 2023) to allow users to exclude undesired concepts or elements from the generated output. Negative prompts are incorporated through classifier-free guidance during the decoding process (Ho & Salimans, 2021). For example, the official prompt guide from Playground suggests using phrases like “ugly, deformed hands” to discourage unwanted aesthetics. [^4] https://playground.com/prompt-guide/negative-prompts Despite their utility, negative prompts are currently under-studied as a means to exclude specific copyrighted elements from generated outputs.
We test negative prompts as a mitigation strategy based on the important anchoring keywords selected via our methods in §3. Specifically, negative prompts are “Copyrighted character” or specific target’s name paired with one of the following options: EmbeddingSim Co-Occurrence LM-Ranked options: 1) $k$ LM-Ranked keywords; 2) $k$ EmbeddingSim keywords; 3) $k$ Co-Occurrence keywords; 4) $k$ EmbeddingSim + $k$ Co-Occurrence keywords.
In addition, we propose to further combine prompt rewriting and negative prompts to strengthen the copyright protection.
5 Experiments and Discussion
@@
To ensure a clear understanding and better control over model behaviors, our evaluation primarily focuses on four state-of-the-art open-source image generation models: Playground v2.5 (Li et al., 2024a), Stable Diffusion XL (SDXL) (Podell et al., 2024), PixArt-$\alpha$ (Chen et al., 2024), and DeepFloyd IF (StabilityAI, 2023), DALL·E 3 Betker et al. (2023), as well as one video generation model, VideoFusion (Luo et al., 2023). [^5] Video generation pipelines can be broadly categorized into the two types: 1) image generation model followed by an image-to-video model, and 2) a direct text-to-video pipeline. For models in the first category (eg. Stable Video Diffusion (Blattmann et al., 2023)), our findings on image generation models are also applicable. Therefore, we focus our video experiments on models of the second type, e.g. VideoFusion (Luo et al., 2023). The configuration details for each model used in our experiments can be found in Section C.2.
Our main analysis focuses on the Playground v2.5 due to its superior generation quality. We also report results for other models in Section F.5.
@@ 5.1 Identifying Prompts That Generate Copyrighted Characters
First, not too surprisingly, we have verified that when using character names, ~60% of tested characters can be generated.  [^6] However, we find that models are not robust to misspellings of character names and generally do not result in generation of characters even with minor misspellings, see Section F.4. For the remainder of this section, we focus on indirect anchoring, where the prompt does not explicitly contain the character’s name. We examine the effect of two types of indirect anchors: textual descriptions and keywords, as well as how to automatically discover them (§3), by checking $\mathsf{DETECT}$, the number of detected copyrighted characters in the generation.
@@
Figure 3: Selection of generated images by Playground v2.5 that are detected as the requested characters by the GPT-4V evaluator. As shown, the model is able to generate images that look highly similar to the required character with (a) or w/o the character’s name in the prompt (b, c).
60-word descriptions lead to the generation of ~48% characters. As described in §3, the first type of indirect prompt uses around 60 words to describe a character’s visual appearance. Despite omitting character names, these descriptions often lead to successful character generation, as shown in Figure 2. Furthermore, prompts with higher embedding similarity to a character’s name tend to generate that character more reliably. Among 100 randomly generated 60-word descriptions per character, the top-ranked description by embedding similarity generates 24 characters successfully, versus only 16 for the bottom-ranked (see Section F.3).
We note that a concurrent study by Kim et al. (2024) also examine keywords potentially important for image generation, but only include the character name along with the associated movie or TV program as keywords. They also show that LLM-optimized descriptions can generate images similar to copyrighted characters on proprietary models such as ChatGPT, Copilot, and Gemini. However, their optimized prompts do not explicitly exclude the characters’ names.
A few keywords, especially those with most frequent co-occurrence with character names in LAION, also easily generate copyrighted characters. We examine the effectiveness of keywords with top co-occurrence frequency with the copyrighted characters’ names (§3) and visualize results in Figure 2. For most selections, we find that keywords chosen from LAION are more effective than using other methods. This is likely because this multimodal dataset is more common in training of image generation models compared to the other text-only ones. Notably, using 5 LAION keywords can almost match performance of using 60-word descriptions. Top 20 LAION and embedding-ranked keywords can both generate more copyrighted characters than using the more detailed paragraph descriptions. Figure 3 shows some examples of these generated images with the descriptions and keywords discussed above.
Descriptions and identified keywords also transfer to generating characters from DALL·E 3 and video models. We further test indirect anchors on production-level models, such as DALL·E 3.
@@
Surprisingly, indirect anchors like descriptions can still bypass system safeguards and result in the generation of copyrighted characters (Figure 4). This further suggests that current safeguards are not fully effective. More results can be found in Appendix E.In addition, we also test indirect anchors on the video generation model VideoFusion (Luo et al., 2023)(see Figure 4 for examples). We compare selection methods for indirect anchors in Figure 2. LAION is the most useful corpus for identifying such keywords, and has a smaller gap to 60-word description on video generation compared to image generation.
@@ 5.2 Mitigation Effectiveness
The next question is: can we effectively prevent the models from recreating these copyrighted characters? We mainly evaluate the intervention strategies discussed in §4, specifically: 1) using prompt rewriting only, 2) using negative prompts only, [^7] To effectively apply the proposed negative prompts, model deployers need a mechanism to detect the identity of the intended copyrighted character (if any) from the user’s prompt. As the primary focus of this work is not end-to-end system building but the evaluation of specific mitigation methods, we assume the existence of such a method. However, we provide more discussion on this in Section F.2 and demonstrate two possible implementations for detecting whether a prompt may reference (directly or indirectly) a popular character. and 3) combining negative prompts and prompt rewriting.
We evaluate these strategies on CopyCat using $\mathsf{DETECT}$ and $\mathsf{CONS}$ as described in §2.2: $\mathsf{DETECT}$ counts the number of detected copyrighted characters, and $\mathsf{CONS}$ measures the image’s consistency with user input. A good mitigation strategy achieves low $\mathsf{DETECT}$ and high $\mathsf{CONS}$. We run each strategy three times and report the mean and standard deviation of $\mathsf{DETECT}$ and $\mathsf{CONS}$ in Table 1.
Prompt rewriting alone is not entirely effective at eliminating outputs similar to copyrighted characters. Our evaluation starts with prompt rewriting (§4), which has been adopted as an intervention strategy for production-level models like DALL·E. However, as demonstrated in Table 1, solely adopting prompt rewriting can only reduce $\mathsf{DETECT}$ from 30 to 14. Nonetheless, an advantage of prompt rewriting is that the $\mathsf{CONS}$ scores modestly improve, likely due to the rewritten prompts containing more detailed information.
@@
Table 2: The combination of prompt-rewriting and negative prompts (target’s name & 5 EmbeddingSim & 5 Co-Occurrence-LAION keywords) can significantly reduce $\mathsf{DETECT}$ while mostly preserving $\mathsf{CONS}$ across all 5 open-source models tested, making it a promising candidate for copyright-protection intervention.
@@
We then investigate potential reasons for prompt rewriting sometimes failing. Specifically, we calculate the average number of Top-5 LAION keywords present in rewritten prompts that result in $\mathsf{DETECT}=0$ (success) and $\mathsf{DETECT}=1$ (failure). We find that the failed rewritten prompts contain on average more LAION keywords—$0.667$ for failure cases compared to $0.387$ for success cases. Similarly, we also observe that failing rewritten prompts tend to share higher embedding similarity with the character’s name (see Section F.3). This again suggests the existence of indirect anchors, and potentially their inclusion in rewritten prompts could impair this strategy.
@@
Using negative prompts improves elimination of similar output with modest impact on consistency. In addition to existing countermeasures like prompt-rewriting, we also explore negative prompts (§4). Specifically, we use keywords identified with different methods (§4, with $k=5$) as negative prompts. We generally observe that including Co-Occurrence-LAION results in higher reduction in $\mathsf{DETECT}$ compared to including LM-Ranked and EmbeddingSim (Table 1). This shows that LAION keywords are more effective than those ranked by LM or embedding space distance.
@@
Including character names in the negative prompt is also helpful. As shown in Table 1, compared to the upper half, the lower half (target name included in negative prompt) consistently has lower $\mathsf{DETECT}$ scores. [^8] We also examine the effectiveness of adding the character name to the negative prompt when user input does not contain the character name and also find a consistent effect. For example, in the case of paragraph-length descriptions, the number of detected characters is reduced by over $50\%$ while maintaining consistency (see Section F.1). Incorporating LAION keywords into the negative prompts in addition to character name further reduces $\mathsf{DETECT}$. The combination of these words in the negative prompt significantly reduces the original $\mathsf{DETECT}$ score from 30 to 4. Notably, the addition of negative prompts does not significantly impair generated image’s consistency with user’s intended prompt, as the $\mathsf{CONS}$ scores typically remain similar or only slightly lower compared to the no intervention setting, but still substantially above $0.33$, the value which indicates very high consistency (see Section C.4). Figure 5 and Figure 12 (in Section F.4) visualize some qualitative examples.
@@
Combining prompt rewriting and negative prompts shows promise for elimination of similar output. Finally, we combine prompt rewriting and negative prompts. Specifically, we send the rewritten prompts as inputs to the image generation models. Then we apply negative prompts during generation. Surprisingly, as demonstrated in Table 2, this simple technique is already quite promising in alleviating copyright concerns and is effective across all open-source models evaluated. [^9] DALL·E does not allow customizing negative prompts. The number of detected copyrighted characters is significantly reduced for all models. Notably, the number of detection decreases to only 5% of the original in the case of DeepFloyd. At the same time, the $\mathsf{CONS}$ scores remain mostly stable. This suggests that despite the pressing concern of image generation models generating copyrighted characters, we can use this simple yet effective method for meaningful mitigation. Figure 5 and Figure 13 (in Section F.4) present some examples. As shown, most generated images still align with the user’s intent in the sense that the generated figure is of a similar entity as the requested copyrighted character, but the generation result is already drastically different from the requested copyrighted characters.
@@
Nonetheless, even this combination of strategies is not perfect at stopping the generation of copyrighted characters, which calls for more future research efforts.
@@ 6 Related Work
Diffusion models is a type of generative models that synthesize images through two intertwined processes: the forward diffusion and the reverse diffusion paths (Rombach et al., 2021; Podell et al., 2023). In the forward diffusion process, an image gradually transitions from its original state to a fully noised version by incrementally adding noise. The reverse process aims to reconstruct the original image from this noisy state. These models can approach the reverse process in two ways: by either predicting the clean image directly at each step or by estimating the noise to be subtracted from the noisy image. Training diffusion models requires extensive datasets, such as LAION-5B, which consists of a vast collection of publicly accessible copyrighted materials (Schuhmann et al., 2022). As these models evolve, diffusion models can generate copies of samples from their training data (Carlini et al., 2023b; Vyas et al., 2023), which raises potential concerns regarding privacy and copyright. Recent works have explored some potential pathways to suppress certain concepts from being generated in the diffusion process (Kumari et al., 2023; Li et al., 2024b). While these methods further fine-tune models and address memorized styles and images individually, we aim to examine operationalizable ways to add copyright protection without updating the parameters.
Copyright and Generative Models
@@
Recent studies have delved into the copyright implications of generative models such as diffusion models and language models (Sag, 2018; Henderson et al., 2023; Lee et al., 2024; Sag, 2023). Lee et al. (2024), Sag (2023), and Henderson et al. (2023) in particular point to copyrighted characters as a challenging legal area. Each of them note that it may be possible for characters to be generated even when users don’t explicitly input the character name, though they do not systematically evaluate this phenomenon.
@@
Others have demonstrated that these models can potentially reconstruct or replicate copyrighted content from their training data (Carlini et al., 2020, 2023b). Efforts to mitigate these risks include provable copyright protection strategies inspired by differential privacy (Vyas et al., 2023), decoding-time prevention (Golatkar et al., 2024) that guide the generation process away from copyright concepts and model editing and unlearning that aim to remove copyrighted content from model weights (Chefer et al., 2023; Zhang et al., 2023). However, the legal framework remains underdeveloped, posing a significant challenge as the capabilities of these generative models continue to advance. A concurrent study by Kim et al. (2024) leverages a large language model optimizer to generate prompts that potentially maximize the likelihood of generating copyrighted content in proprietary image-generation models. Similarly, Zhang et al. (2024) focus on building attacks that can generated particular concepts—where concepts are broadly defined and include some copyrighted characters. These works focus on attacks and do not explore effective mitigation methods. Our work, however, focuses on building an evaluation framework for understanding the effectiveness of defenses. This necessarily includes similarly identifying indirect anchors that might generate copyrighted characters. These studies also focus on longer prompts, while our work demonstrates the possibility of such violations with just a few keywords (around five).
7 Limitations and Future Work

### v2

@@ 4 Mitigation Strategies
Prompt rewriting is an existing mitigation used in production-level systems such as DALL·E. Specifically, the DALL·E interface contains a prompt-rewriting step that first processes the user’s text input into a format that DALL·E can use to generate images and comply with OpenAI’s policies, such as avoiding copyrighted content. Prompt rewriting approaches currently adopted by companies reflect our evaluation framework: they encourage policy compliance (e.g., avoiding the generation of specific copyrighted characters like Mario) while staying close to the user’s intent (e.g., if the user requests a plumber, still generating a plumber). In order to simulate the prompt-rewriting pipeline, we query GPT-4 with the DALL·E’s system prompt obtained via prompt extraction methods (see full template and details in § C.6) and the keywords or descriptions to be rewritten.
Example of applying prompt rewriting for ‘Mario’ ‘Create an image of a fictional character inspired by the world of classic video games. He is a middle-aged man of Italian descent, with a robust physique, and typically clad in a red shirt and blue overalls. His most distinctive features include a bushy mustache and a red cap… ’
@@
Negative prompts are often used in deployed diffusion model deployments (Playground AI, 2023) to allow users to exclude undesired concepts or elements from the generated output. Negative prompts are incorporated through classifier-free guidance during the decoding process (Ho & Salimans, 2021). For example, the official prompt guide from Playground suggests using phrases like “ugly, deformed hands” to discourage unwanted aesthetics.https://playground.com/prompt-guide/negative-prompts Despite their utility, negative prompts are currently under-studied as a means to exclude specific copyrighted elements from generated outputs.
We test negative prompts as a mitigation strategy based on the important anchoring keywords selected via our methods in § 3. Specifically, negative prompts are “Copyrighted character” or specific target’s name paired with one of the following options: 1) $k$ LM-Ranked keywords; 2) $k$ EmbeddingSim keywords; 3) $k$ Co-Occurrence keywords; 4) $k$ EmbeddingSim + $k$ Co-Occurrence keywords. In addition,further combine prompt rewriting and negative prompts to reduce the likelihood of generating copyright characters.
5 Experiments and Analysis
@@
To ensure a clear understanding and better control over model behaviors, our evaluation primarily focuses on four state-of-the-art open-source image generation models: Playground v2.5 (Li et al., 2024a), Stable Diffusion XL (SDXL) (Podell et al., 2024), PixArt-$\alpha$ (Chen et al., 2024), and DeepFloyd IF (StabilityAI, 2023), DALL·E 3 Betker et al. (2023), as well as one video generation model, VideoFusion (Luo et al., 2023).§§§Video generation pipelines can be broadly categorized into the two types: 1) image generation model followed by an image-to-video model, and 2) a direct text-to-video pipeline. For models in the first category (eg. Stable Video Diffusion (Blattmann et al., 2023)), our findings on image generation models are also applicable. Therefore, we focus our video experiments on models of the second type, e.g. VideoFusion (Luo et al., 2023). The configuration details for each model used in our experiments can be found in § C.2. Our main analysis focuses on the Playground v2.5 due to its superior generation quality. We also report results for other models in § E.5.
@@ 5.1 Identifying Prompts That Generate Copyrighted Characters
Figure 2: Selection of generated images by Playground v2.5 that GPT-4V detects as the requested characters. As shown, the model is able to generate images that look highly similar to the required character with (a) or w/o the character’s name in the prompt (b, c).
First, not too surprisingly, we have verified that when using character names, ~60% of tested characters can be generated, also highlighted in the upper-left of Tb. 1.However, we find that models are not robust to misspellings of character names and generally do not result in generation of characters even with minor misspellings, see § E.4. For the remainder of this section, we focus on indirect anchoring, where the prompt does not explicitly contain the character’s name. We examine the effect of two types of indirect anchors: textual descriptions and keywords, as well as how to automatically discover them (§ 3), by checking $\mathsf{DETECT}$, the number of detected copyrighted characters in the generation.
60-word descriptions lead to the generation of ~48% characters. Despite omitting character names, these descriptions often lead to successful character generation, as shown in the horizontal dashed line in Fig. 3. Furthermore, prompts with higher embedding similarity to a character’s name tend to generate that character more reliably (see § E.3). Among 100 randomly generated 60-word descriptions per character, the top-ranked description by embedding similarity generates 24 characters successfully, versus only 16 for the bottom-ranked (see § E.3).
A few keywords, especially those with most frequent co-occurrence with character names in LAION, also easily generate copyrighted characters. We examine the effectiveness of keywords with top co-occurrence frequency with the copyrighted characters’ names (§ 3) and visualize results in Fig. 3. For most selections, we find that keywords chosen from LAION are more effective than using other methods. This is likely because this multimodal dataset is more common in training of image generation models compared to the other text-only ones. Notably, using 5 LAION keywords can almost match performance of using 60-word descriptions. Top 20 LAION and embedding-ranked keywords can both generate more copyrighted characters than using the more detailed paragraph descriptions. Fig. 2 shows some examples of these generated images with the descriptions and keywords discussed above.
@@
Descriptions and identified keywords also transfer to generating characters from DALL·E 3 and video models. We further test indirect anchors on production-level models, such as DALL·E 3. Surprisingly, some indirect anchors like descriptions can still bypass system safeguards and result in the generation of copyrighted characters (Fig. 4). This further suggests that current safeguards are not fully effective. More results can be found in § D. In addition, we also test indirect anchors on the video generation model VideoFusion (Luo et al., 2023) (see Fig. 4 for examples). We compare selection methods for indirect anchors in Fig. 3. LAION is the most useful corpus for identifying such keywords, and has a smaller gap to 60-word description on video generation compared to image generation.
@@ 5.2 Mitigation Effectiveness
The next question is: can we effectively prevent the models from recreating these copyrighted characters? We mainly evaluate the intervention strategies discussed in § 4, specifically: 1) using prompt rewriting only, 2) using negative prompts only,To effectively apply the proposed negative prompts, model deployers need a mechanism to detect the identity of the intended copyrighted character (if any) from the user’s prompt. As the primary focus of this work is not end-to-end system building but the evaluation of specific mitigation methods, we assume the existence of such a method. However, we provide more discussion on this in § E.2 and demonstrate two possible implementations for detecting whether a prompt may reference (directly or indirectly) a popular character. and 3) combining negative prompts and prompt rewriting.
We evaluate these strategies using the evaluation framework as described in § 2: $\mathsf{DETECT}$ counts the number of detected copyrighted characters, and $\mathsf{CONS}$ measures the image’s consistency with user input. A good mitigation strategy achieves low $\mathsf{DETECT}$ and high $\mathsf{CONS}$. We run each strategy three times and report the mean and standard deviation of $\mathsf{DETECT}$ and $\mathsf{CONS}$ in Tb. 1.
Prompt rewriting alone is not entirely effective at eliminating outputs similar to copyrighted characters. Our evaluation starts with simulating prompt rewriting (§ 4), which has been adopted as an intervention strategy for production-level models like DALL·E. However, as demonstrated in Tb. 1, solely adopting prompt rewriting can only reduce $\mathsf{DETECT}$ from 30 to 14. Nonetheless, an advantage of prompt rewriting is that the $\mathsf{CONS}$ scores modestly improve, likely due to the rewritten prompts containing more detailed information.
@@
Table 2: The combination of prompt-rewriting and negative prompts (target’s name & 5 EmbeddingSim & 5 Co-Occurrence-LAION keywords, shown on the right half of the table) can significantly reduce $\mathsf{DETECT}$ while mostly preserving $\mathsf{CONS}$ compared to no-intervention baseline (left half of the table) across all 5 open-source models tested, making it a promising candidate for copyright-protection intervention.
@@
We then investigate potential reasons for prompt rewriting sometimes failing. Specifically, we calculate the average number of Top-5 LAION keywords present in rewritten prompts that result in $\mathsf{DETECT}=0$ (success) and $\mathsf{DETECT}=1$ (failure). We find that the failed rewritten prompts contain on average more LAION keywords—$0.667$ for failure cases compared to $0.387$ for success cases. Similarly, we also observe that failing rewritten prompts tend to share higher embedding similarity with the character’s name (see § E.3). This again suggests the existence of indirect anchors, and potentially their inclusion in rewritten prompts could impair this strategy.
@@
Using negative prompts improves elimination of similar output with modest impact on consistency. In addition to existing countermeasures like prompt-rewriting, we also explore negative prompts (§ 4). Specifically, we use keywords identified with different methods (§ 4, with $k=5$) as negative prompts. We generally observe that including Co-Occurrence-LAION results in higher reduction in $\mathsf{DETECT}$ compared to including LM-Ranked and EmbeddingSim (Tb. 1).
@@
Including character names in the negative prompt is also helpful. As shown in Tb. 1, compared to the upper half, the lower half (target name included in negative prompt) consistently has lower $\mathsf{DETECT}$ scores.******We also examine the effectiveness of adding the character name to the negative prompt when user input does not contain the character name and also find a consistent effect. In the case of paragraph-length descriptions, the number of detected characters is reduced by over $50\%$ while maintaining consistency (see § E.1). Incorporating LAION keywords into the negative prompts in addition to character name further reduces $\mathsf{DETECT}$. The combination of these words in the negative prompt significantly reduces the original $\mathsf{DETECT}$ score from 30 to 4. Notably, the addition of negative prompts does not significantly impair generated image’s consistency with user’s intended prompt, as the $\mathsf{CONS}$ scores typically remain similar or only slightly lower compared to the no intervention setting, but still substantially above $0.33$, the value which indicates very high consistency (see § C.4). Fig. 5 and Fig. 12 (in § E.4) visualize some qualitative examples.
@@
Combining prompt rewriting and negative prompts shows promise for elimination of similar output. Finally, we combine prompt rewriting and negative prompts. Specifically, we send the rewritten prompts as inputs to the image generation models. Then we apply negative prompts during generation. Surprisingly, as demonstrated in Tb. 2, this simple technique is already quite promising in alleviating copyright concerns and is effective across all open-source models evaluated.††††††DALL·E does not allow customizing negative prompts. The number of detected copyrighted characters is significantly reduced for all models. Notably, the number of detection decreases to only 5% of the original in the case of DeepFloyd. At the same time, the $\mathsf{CONS}$ scores remain mostly stable. This suggests that despite the pressing concern of image generation models generating copyrighted characters, we can use this simple yet effective method for meaningful mitigation. Fig. 5 and Fig. 13 (in § E.4) present some examples. As shown, most generated images still align with the user’s intent, keeping key characteristics as the requested copyrighted character, but the generation result is already drastically different from the requested copyrighted characters.
Table 3: Our intervention strategy combining rewriting with negative prompts has a statistically significant effect on $\mathsf{DETECT}$ scores. The negative prompt set refers to Target’s name + 5 EmbeddingSim & 5 Co-Occurrence-LAION keywords. T-values represent one-tailed tests ($\alpha$ = 0.05) and ∗∗ denotes $p<0.05$.
Prompt Mean (Baseline) Mean (with negative prompt set) T-value
$\pm$ SD $\pm$ SD
Target’s Name $30.33\pm 1.89$ $20.67\pm 0.47$ $8.59^{**}$
Rewritten Prompt $14.33\pm 2.62$ $4.33\pm 0.47$ $6.51^{**}$
@@
We show statistical significance of our highlighted intervention’s effect on reducing $\mathsf{DETECT}$ in Tb. 3. Nonetheless, even this combination of strategies is not perfect at stopping the generation of copyrighted characters, which calls for more future research efforts.
@@ 6 Related Work
Diffusion models is a type of generative models that synthesize images through two intertwined processes: the forward and the reverse diffusion paths (Rombach et al., 2021; Podell et al., 2023). In the forward diffusion process, an image gradually transitions from its original state to a fully noised version by incrementally adding noise. The reverse process aims to reconstruct the original image from this noisy state. These models can approach the reverse process in two ways: by either predicting the clean image directly at each step or by estimating the noise to be subtracted from the noisy image. Training diffusion models requires extensive datasets, such as LAION-5B, which consists of a vast collection of publicly accessible copyrighted materials (Schuhmann et al., 2022). As these models evolve, diffusion models can generate copies of samples from their training data (car, 2023; Vyas et al., 2023), which raises potential concerns regarding privacy and copyright. Recent works have explored some potential pathways to suppress certain concepts from being generated in the diffusion process (Kumari et al., 2023; Li et al., 2024b). While these methods further fine-tune models and address memorized styles and images individually, we aim to examine operationalizable ways to add copyright protection without updating the parameters.
Copyright and generative models.
@@
Recent studies have delved into the copyright implications of generative models such as diffusion models and language models (Sag, 2018; Henderson et al., 2023; Lee et al., 2024; Sag, 2023; Min et al., 2023; Shi et al., 2024). Lee et al. (2024), Sag (2023), and Henderson et al. (2023) in particular point to copyrighted characters as a challenging legal area. Each of them note that it may be possible for characters to be generated even when users don’t explicitly input the character name, though they do not systematically evaluate this phenomenon.
@@
Others have demonstrated that these models can potentially reconstruct or replicate copyrighted content from their training data (Carlini et al., 2020; car, 2023). Efforts to mitigate these risks include provable copyright protection strategies inspired by differential privacy (Vyas et al., 2023), decoding-time prevention (Golatkar et al., 2024) that guide the generation process away from copyright concepts and model editing and unlearning that aim to remove copyrighted content from model weights (Gong et al., 2024; Chefer et al., 2023; Zhang et al., 2023). Ma et al. (2024) introduces benchmark for measuring copyright infringement unlearning from text-to-image diffusion models. Another recent study by Kim et al. (2024) also examines keywords potentially important for image generation, but only includes the character name along with the associated movie or TV program as keywords. They also show that LLM-optimized descriptions can generate images similar to copyrighted characters on proprietary models such as ChatGPT, Copilot, and Gemini. However, their optimized prompts do not explicitly exclude the characters’ names. Similarly, Zhang et al. (2024a) focuses on building attacks that can generated particular concepts,including some copyrighted characters. While these works focus on attacks and do not explore effective mitigation methods, our work focuses on building an analysis framework motivated by legal considerations like substantial similarity test for copyrighted characters and provide more in-depth understanding for both how easy it is to generate copyrighted characters as well as the effectiveness of defenses.
7 Discussions and Limitations
Our work provides an initial step forward for systematically evaluating the likelihood of generating copyrighted characters and the effectiveness of inference-time mitigation strategies. While our LLM-as-a-judge method can make mistakes in judging character similarity (§ C.3), the relatively low error rate does not affect the ranking of our methods nor our main takeaways: Current mitigation strategies are imperfect, and there are still many instances where characters are generated verbatim. There is still plenty of room for improving mitigations.
Note that whether the proposed strategies will be considered sufficient is also a jurisdiction-dependent question. There are limits to the proposed mitigation strategies and current evaluation, and the metrics we propose should not be treated as a universal optimization target. Nevertheless, our analysis suggests that current mitigations are lacking and more should be done to assess compliance with one’s policy. We hope that this work can help understand the gaps in these strategies and facilitate more complete assessment of model and system design in future research.